(3p)^4(3p)^-7 I'm supposed to simplify and can't seem to figure this out. I distributed the exponent 4 on the first term and got 3^4 p^4, don't know what to do from here. thank you!
Answer by sdmmadam@yahoo.com(530)

(3p)^4(3p)^-7 I'm supposed to simplify and can't seem to figure this out. I distributed the exponent 4 on the first term and got 3^4 p^4, don't know what to do from here. thank you!
(3p)^4(3p)^-7
= [(3)^4 X (p)^4] X [(3)^(-7) X (p)^(-7)]
[Using (ab)^m = (a^m)X(b^m) ]
=(3)^(4-7)X (p)(4-7)
[Using (a)^m X (a)^n = (a)^(m+n) and here in the problem n is negativae]
=(3)^(-3)X (p)^(-3)
=(3p)^-3
=1/[(3p)^(+3)] (using {(a)^(-m) = 1/[(a)^(+m)]} )
=1/[3^3p^3]
=1/(27p^3)